OpenStudy (anonymous):

I understand the idea of implicit differentiation, but actually working it out doesn't quite work out for me. Here's an example. 100x^.75y^.75= 135540 I do the product rule first and end up with 100x^.75(.25y^-.75dy/dx) + y^.25(75x^-.25)=0 i understand I have to get dy/dx by itself, but i dont see how they got 25x^.75/y^.75=-75y^.25/x^.25 Can you help?

OpenStudy (anonymous):

25x^.75/y^.75=-75y^.25/x^.25 ?

OpenStudy (anonymous):

i'll rewrite it

OpenStudy (anonymous):

\[\frac{ d }{dx } 100(xy)^{\frac{ 3 }{4 }} =\frac{ 3 }{4 }100(xy)^{\frac{ -1 }{ 4 }}(y+xy \prime )\]

OpenStudy (anonymous):

right but i dont see how they got to this part\[\frac{ 25x ^{.75} }{ y ^{.75} }* dy/dx \] = \[\frac{ -75y ^{.75} }{ x ^{.25} } \]

OpenStudy (anonymous):

you sure you wrote the original correctly?

OpenStudy (anonymous):

it's 100x\[100x ^{3/4}y ^{1/2} = 135,540\] it's an implicit differentiation problem

OpenStudy (anonymous):

..

OpenStudy (anonymous):

100x^.75y^.75= 135540 not the same...

OpenStudy (anonymous):

o didnt relaize i wrote .75 instead of .25.

OpenStudy (anonymous):

that's not the problem actually. I wouldn't recommend converting to decimals. you might be causing yourself confusion.

OpenStudy (anonymous):

they gave it to me in decimal form, isnt .75=3/4

OpenStudy (anonymous):

yes but 1/2 isn't .25

OpenStudy (anonymous):

or .75

OpenStudy (anonymous):

im soo gone right now lol. ur right .25=1/4

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

100x^(3/4) y^(1/2)=135,540 so is that your problem?

OpenStudy (anonymous):

1/4 instead of 1/2

OpenStudy (anonymous):

ah

OpenStudy (anonymous):

whoops dy/dx = -3y/x?

OpenStudy (anonymous):

yes it's -3y/x. can you explain how you got there.

OpenStudy (anonymous):

sure:

OpenStudy (anonymous):

ignore the 100, product rule: \[\frac{ d }{dx } ( x ^{\frac{ 3 }{4 }}y ^{\frac{ 1 }{ 4 }} = C)\]

OpenStudy (anonymous):

ok so treat it as a constant

OpenStudy (anonymous):

\[\huge \frac{ 3 }{4 }x ^{\frac{ -1 }{ 4 }}y ^{\frac{ 1 }{ 4 }} + x ^{\frac{ 3 }{ 4 }}\frac{ 1 }{4 }y ^{\frac{ -3 }{ 4}}y \prime =0\]

OpenStudy (anonymous):

\[\huge x ^{\frac{ 3 }{ 4 }}\frac{ 1 }{ 4 }y ^{\frac{ -3 }{ 4 }}y \prime = -\frac{ 3 }{ 4}x ^{\frac{ -1 }{ 4 }}y ^{\frac{ 1 }{ 4 }}\]

OpenStudy (anonymous):

ok i'm following.

OpenStudy (anonymous):

\[\huge y \prime = \frac{ -\frac{ 3 }{ 4}x ^{\frac{ -1 }{ 4 }}y ^{\frac{ 1 }{ 4 }} }{ x ^{\frac{ 3 }{ 4 }}\frac{ 1 }{ 4 }y ^{\frac{ -3 }{ 4 }} }\]

OpenStudy (anonymous):

make all exponents positive. divide -3/4 by 1/4

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

im working it out right now, hold on

OpenStudy (anonymous):

ok i see where the 3 comes from. now i have \[\frac{ y ^{3/4} y ^{1/4} }{ x ^{3/4}x ^{1/4} }\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ie y/x

OpenStudy (anonymous):

how do those cancel out? Is it because of the same exponent regardless of the base?

OpenStudy (anonymous):

nevermind i got it. rules of exponent. thank you for your help

OpenStudy (anonymous):

sure.